### v1

# mbrs: A Library for Minimum Bayes Risk Decoding

Abstract
Minimum Bayes risk (MBR) decoding is a decision rule of text generation tasks that outperforms conventional maximum a posterior (MAP) decoding using beam search by selecting high-quality outputs based on a utility function rather than those with high-probability. Typically, it finds the most suitable hypothesis from the set of hypotheses under the sampled pseudo-references. mbrs is a library of MBR decoding, which can flexibly combine various metrics, alternative expectation estimations, and algorithmic variants. It is designed with a focus on speed measurement and calling count of code blocks, transparency, reproducibility, and extensibility, which are essential for researchers and developers. We published our mbrs as an MIT-licensed open-source project, and the code is available on GitHub [^1] GitHub: https://github.com/naist-nlp/mbrs,222 YouTube: https://youtu.be/4qeHpg4PTn0 , [^3] Demo: https://mbrs-demo.streamlit.app/
mbrs: A Library for Minimum Bayes Risk Decoding
Hiroyuki Deguchi,  Yusuke Sakai,  Hidetaka Kamigaito,  Taro Watanabe Nara Institute of Science and Technology (NAIST) {deguchi.hiroyuki.db0, sakai.yusuke.sr9, kamigaito.h, taro}@is.naist.jp
1 Introduction
Text generation has now become a central topic in natural language processing owing to the success of language models. A typical text generation model generates the high-probability text using beam search and other search algorithms, which relies on maximum a posteriori (MAP) decoding; however, recent studies have demonstrated that high-probability texts are not always high-quality. This phenomenon is known as beam search curse and the models sometimes generate pathologically broken texts, e.g., empty sequences, n-gram repetitions, and copies of the inputs (Koehn and Knowles, 2017; Ott et al., 2018; Eikema and Aziz, 2020).
Minimum Bayes risk (MBR) decoding addresses the problems of MAP decoding by determining outputs according to the decision rule based on quality or preference rather than probability. Its effectiveness has been confirmed in statistical automatic speech recognition (Goel and Byrne, 2000) and statistical machine translation Kumar and Byrne (2004), and it has been applied to neural text generation, especially neural machine translation (Eikema and Aziz, 2020; Müller and Sennrich, 2021). Although many variants of MBR decoding have been proposed, there is no common library where we can try the latest algorithms and compare them systematically. It is clearly essential for both researchers and developers to establish a library for the further improvement in MBR decoding.
Figure 1: Workflow overview of mbrs.
We introduce mbrs, a library of MBR decoding, which implements various metrics and algorithms with many useful features for comparisons and the development of new methods. Figure 1 shows the workflow overview of mbrs. In the figure, each module, i.e., “Metric” and “Decoder”, is easily extensible. In addition, we carefully designed mbrs to ensure transparency and reproducibility. For instance, the profiler, which automatically measures how much time a code block consumes and how many times it is called, and the typed dataclass, which returns helpful metadata for analyses, are implemented.
Experiments on the WMT’22 En–De general translation task demonstrated that our mbrs can not only systematically compare various metrics and algorithms but also present the runtime statistics and supplementary information of output texts.
2 Background
2.1 Minimum Bayes risk (MBR) decoding
The goal of MBR decoding is to find the high-quality output text from a candidate list for a given input text. We denote input and output texts as sequences of tokens $\bm{x}\in\mathcal{V}^{\ast}_{X}$ and $\bm{y}\in\mathcal{V}^{\ast}_{Y}$, respectively. Note that $\mathcal{V}^{\ast}_{X}$ and $\mathcal{V}^{\ast}_{Y}$ are the Kleene closures of input and output vocabularies, respectively. Since $\bm{y}$ is searched from the output space $\mathcal{Y}\coloneqq\{\bm{y}\mid\bm{y}\in\mathcal{V}^{\ast}_{Y}\}$, which is an infinite set, we usually find the optimal solution with some approximations.
The most commonly used strategy is maximum a posteriori (MAP) decoding using beam search. The solution of MAP decoding $\bm{y}^{\mathrm{MAP}_{\theta}}\in\mathcal{Y}$ is obtained according to its output probability which is calculated by a text generation model $\theta$ as follows:
$\bm{y}^{\mathrm{MAP}_{\theta}}=\operatorname*{argmax}_{\bm{y}\in\mathcal{Y}}p(\bm{y}|\bm{x};\theta).$ (1)
Recent studies have demonstrated that this high-probability sequence $\bm{y}^{\mathrm{MAP}_{\theta}}$ is not always high-quality. Specifically, as the beam size increases, a sequence with higher probability is obtained; however, such sequences hurt translation performance, a phenomenon known as beam search curse. This often results in generating pathological sequences, e.g., empty sequences, n-gram repetitions, and copies of the input sequence (Koehn and Knowles, 2017; Ott et al., 2018; Eikema and Aziz, 2020).
In contrast, minimum Bayes risk (MBR) decoding, an alternative strategy, is known to be effective in avoiding the problems of MAP decoding (Eikema and Aziz, 2020). It finds the high-quality text from the given candidate list $\mathcal{H}\subseteq\mathcal{Y}$ and is formulated as follows:
$\bm{y}^{\mathrm{MBR_{true}}}=\operatorname*{argmax}_{\bm{h}\in\mathcal{H}}\operatorname*{\mathbb{E}}_{\bm{y}\sim\Pr(\bm{y}|\bm{x})}\left[u(\bm{h},\bm{y})\right],$ (2)
where $\Pr(\bm{y}|\bm{x})$ is the true probability of $\bm{y}$ being generated from $\bm{x}$. $u\colon\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is the utility function that measures the quality or preference of generated text under the given reference text, which is defined as $\bm{h}\succeq\bm{h}^{\prime}\iff u(\bm{h},\bm{r})\geq u(\bm{h}^{\prime},\bm{r})$ where $\succeq$ denotes the preference relation. Since the true probability $\Pr(\bm{y}|\bm{x})$ is unknown, it is replaced with the output probability calculated by a text generation model $\theta$:
$\bm{y}^{\mathrm{MBR}_{\theta}}=\operatorname*{argmax}_{\bm{h}\in\mathcal{H}}\operatorname*{\mathbb{E}}_{\bm{y}\sim p(\bm{y}|\bm{x};\theta)}\left[u(\bm{h},\bm{y})\right].$ (3)
Because enumerating all possible output texts is infeasible, the expected scores calculated by the utility function (expected utility; EU) are estimated using pseudo-references that are sampled according to the output probabilities. Let $\hat{\mathcal{R}}$ be a bag (a.k.a multiset) of sampled pseudo-references and $\mathcal{R}$ is the support set of $\hat{\mathcal{R}}$, i.e., a set of distinct elements of $\hat{\mathcal{R}}$, they are formulated as follows:
$\hat{\mathcal{R}}$$\coloneqq\{\bm{y}_{i}\in\mathcal{Y}\mid\bm{y}_{i}\sim p(\bm{y}|\bm{x};\theta)\}_{i=1}^{|\hat{\mathcal{R}}|},$ (4)
$\mathcal{R}$ $\coloneqq\operatorname{Supp}(\hat{\mathcal{R}})=\{\bm{y}\in\mathcal{Y}\mid m_{\hat{\mathcal{R}}}(\bm{y})>0\},$ (5)
[…]
$p_{\mathrm{MC}}(\bm{r}|\bm{x};\hat{\mathcal{R}})$ $\coloneqq\frac{m_{\hat{\mathcal{R}}}(\bm{r})}{|\hat{\mathcal{R}}|},$ (6)
$\mu_{\mathrm{MC}}(\bm{h};\hat{\mathcal{R}})$ $\coloneqq\sum_{\bm{r}\in\operatorname{Supp}(\hat{\mathcal{R}})}p_{\mathrm{MC}}(\bm{r}|\bm{x};\hat{\mathcal{R}})u(\bm{h},\bm{r}),$ (7)
$\bm{y}^{\mathrm{MBR}_{\theta}^{\mathrm{MC}}}$ $=\operatorname*{argmax}_{\bm{h}\in\mathcal{H}}\mu_{\mathrm{MC}}(\bm{h};\hat{\mathcal{R}}),$ (8)
or using the model-based method (Jinnai et al., 2024) based on the output probability:
$\mu_{\mathrm{MB}}(\bm{h};\mathcal{R},\theta)$ $\coloneqq\sum_{\bm{r}\in\mathcal{R}}p_{\mathrm{MB}}(\bm{r}|\bm{x};\mathcal{R},\theta)u(\bm{h},\bm{r}),$ (9)
$\bm{y}^{\mathrm{MBR}_{\theta}^{\mathrm{MB}}}$ $=\operatorname*{argmax}_{\bm{h}\in\mathcal{H}}\mu_{\mathrm{MB}}(\bm{h};\mathcal{R},\theta),$ (10)
where $p_{\mathrm{MB}}$ is the normalized output probability over a set of pseudo-references, as follows:
$p_{\mathrm{MB}}(\bm{r}|\bm{x};\mathcal{R},\theta)\coloneqq\frac{p(\bm{r}|\bm{x};\theta)}{\sum_{\bm{r}^{\prime}\in\mathcal{R}}p(\bm{r}^{\prime}|\bm{x};\theta)}.$ (11)
Typically, the hypotheses themselves are regarded as the pseudo-references $\hat{\mathcal{R}}$ or $\mathcal{R}$.
2.2 Variations in MBR decoding algorithms
While MBR decoding can generate higher-quality texts compared with MAP decoding, the EU estimation requires the time complexity of $\mathcal{O}(|\mathcal{H}||\hat{\mathcal{R}}|)$ or $\mathcal{O}(|\mathcal{H}||\mathcal{R}|)$, which is time-consuming. To address the issue, efficient variants of MBR decoding have been proposed. One is the reference aggregation, which approximates the EU using an aggregated representation of pseudo-references in the feature space (RAMBR) (DeNero et al., 2009; Vamvas and Sennrich, 2024). Similarly, Deguchi et al. (2024) cluster pseudo-references in the embedding space and use the centroid representations for the EU estimation (CBMBR). Another approach is hypothesis pruning, which prunes hypotheses that are not likely to be selected (PruneMBR) (Cheng and Vlachos, 2023). Trabelsi et al. (2024) proposed probabilistic MBR (PMBR) that reduces the number of the utility function calls. PMBR first calculates the utility scores using only sampled pairs of a hypothesis and reference instead of all pairs. Then, it approximately computes the hypothesis–pseudo-reference utility score matrix using the low-rank matrix completion (Zachariah et al., 2012).
2.3 Problems of existing implementations
Currently, MBR decoding has drawn attention from research communities, and although various methods have been proposed, there is no common library that includes many of the latest studies, making it hard to compare the quality and speed systematically. mbr-nmt555mbr-nmt: https://github.com/roxot/mbr-nmt is the original implementation of Monte Carlo estimation (Eikema and Aziz, 2022), but it does not support other later methods or metrics. COMET is known as an evaluation metric, but its framework includes a command for MBR decoding, comet-mbr (Rei et al., 2020; Fernandes et al., 2022). It only supports COMET as an evaluation metric and the Monte Carlo estimation as a decoding method. Vamvas and Sennrich (2024) released mbr666mbr: https://github.com/ZurichNLP/mbr which is highly integrated into huggingface’s transformers (Wolf et al., 2020), but it only comprises their work and the vanilla MBR decoding, and it cannot be combined with other frameworks like fairseq (Ott et al., 2019) and black-box large language model services.
Now, to further advance MBR decoding — a powerful technique for improving the quality of text generation — a systematic and shared library is clearly essential for both researchers and developers.
3 Our Library: mbrs
Our library mbrs is mainly implemented on Python and PyTorch (Paszke et al., 2019). It finds the most suitable output from the given hypotheses.
3.1 Main components
Metrics
Metrics are the collections of various evaluation metrics. Basically, reference-based metrics are implemented that can be used as utility functions, but reference-free metrics are also implemented for N-best list reranking. The following are the available metrics in our current repository:
  • BLEU (Papineni et al., 2002)
  • Translation Edit Rate (TER) (Snover et al., 2006)
  • chrF (Popović, 2015)
  • COMET (Rei et al., 2020, 2022a)
  • COMETkiwi (Rei et al., 2022b)
  • XCOMET (Guerreiro et al., 2023)
  • BLEURT (Sellam et al., 2020)
Decoders
Decoder Available metrics Time complexity
MBR: Vanilla (Eikema and Aziz, 2020, 2022) all $\mathcal{O}(|\mathcal{H}||\hat{\mathcal{R}}|)$
PruneMBR: Confidence-based pruning (Cheng and Vlachos, 2023) all N/A
RAMBR: Reference aggregation BLEU, chrF, and COMET $\mathcal{O}(|\mathcal{H}|+|\hat{\mathcal{R}}|)$
     on BLEU: Aggregate N-gram counts and length DeNero et al. (2009)
     on chrF: Aggregate N-gram counts Vamvas and Sennrich (2024)
     on COMET: Aggregate sentence embeddings Vamvas and Sennrich (2024); Deguchi et al. (2024)
CBMBR: Centroid-based reference aggregation (Deguchi et al., 2024) COMET $\mathcal{O}(|\mathcal{H}|k+|\hat{\mathcal{R}}|k)$
PMBR: Low-rank matrix completion (Trabelsi et al., 2024) all $\mathcal{O}\left(\frac{|\mathcal{H}||\hat{\mathcal{R}}|}{r}\right)$
Table 1: Implementation list of MBR decoding variants. The time complexity is based on the Monte Carlo estimation. Since PruneMBR dynamically decides the termination condition based on confidence that depends on the input data, we are unable to show the time complexity analytically.
mbrs selects the most suitable output based on MBR decoding or N-best list reranking from the set of hypotheses $\mathcal{H}$. When the hypotheses are reranked using reference-free metrics like COMETkiwi, mbrs performs standard N-best list reranking; otherwise, it performs MBR decoding.
In MBR decoding, we support not only Monte Carlo estimation (Eikema and Aziz, 2022) but also model-based estimation (Jinnai et al., 2024) by passing the log-probabilities of pseudo-references. These estimators can be combined with a variety of MBR decoding algorithms. Table 1 lists the implemented variants of MBR decoding. Note that the aggregation methods in RAMBR depend on each metric. mbrs implements the aggregation methods for the BLEU (DeNero et al., 2009), chrF (Vamvas and Sennrich, 2024), and COMET (Vamvas and Sennrich, 2024; Deguchi et al., 2024) that has been proposed so far [^7] Other metrics, e.g., TER, XCOMET, and BLEURT, cannot aggregate the references due to their calculation (DeNero et al., 2009; Vamvas and Sennrich, 2024; Deguchi et al., 2024).. The listed decoders can be combined with either the Monte Carlo or model-based estimations.
3.2 Interfaces
{CJK}
UTF8ipxg
1from mbrs.metrics import MetricCOMET
2from mbrs.decoders import DecoderMBR
3
4SOURCE = "ありがとう"
5HYPOTHESES = ["Thanks", "Thank you", "Thank you so much", "Thank you.", "thank you"]
6
7metric_cfg = MetricCOMET.Config(
8 model="Unbabel/wmt22-comet-da",
9)
10metric = MetricCOMET(metric_cfg)
11
12decoder_cfg = DecoderMBR.Config()
13decoder = DecoderMBR(decoder_cfg, metric)
14
15output = decoder.decode(HYPOTHESES, references=HYPOTHESES, source=SOURCE, nbest=1)
16
17print(f"Selected index: {output.idx}")
18print(f"Output sentence: {output.sentence}")
19print(f"Expected score: {output.score}")
Listing 1: An example code of MBR decoding using the COMET metric on our mbrs.
mbrs has two interfaces: Python application programming interface (API) and command-line interface (CLI). Listing 1 is an example code via Python API. The detailed references of Python API and CLI are available on Read the Docs [^8] Read the Docs: https://mbrs.readthedocs.io/en/latest/index.html.
3.3 Reproducibility
Fixing random number seed
Some algorithms use random numbers. To ensure the reproducibility of experiments, mbrs generates all random numbers from torch.Generator with a manual seed.
Dataclass/YAML-based configuration
All metrics and decoders are configured using dataclass, making it more robust by using typed variables. In CLI, the command-line arguments are automatically generated and parsed using the configuration dataclasses. Furthermore, instead of specifying arguments directly, the YAML configuration file can also be passed via --config_path option in CLI.
3.4 Transparency
Code block profiler
1$>$ from mbrs import timer
2$>$ from mbrs.metrics import MetricBLEU
3$>$ HYPOTHESES = ["Thank you so much."] * 100
4$>$ metric = MetricBLEU(MetricBLEU.Config())
5$>$ scores = []
6$>$ for h in HYPOTHESES:
7... for r in HYPOTHESES:
8... with timer.measure("score"):
9... scores.append(metric.score(h, r))
10$>$ timer.aggregate().result(nsentences=1)
11[{’name’: ’score’, ’acctime’: 0.6932655478303786, ’acccalls’: 10000, ’ms/call’: 0.06932655478303787, ’ms/sentence’: 693.2655478303786, ’calls/sentence’: 10000.0}]
Listing 2: An example usage of our code block profiler.
One of the key features of mbrs is the code block profiler as shown in Listing 2. It can measure the elapsed time and number of calls within the code block using a context manager, i.e., with statement of Python. After running the program, it automatically aggregates all profilers and reports the statistics. This feature is helpful for identifying the bottleneck of the codes and designing new algorithms.
Metadata analysis
In addition to the configuration, the outputs of the decoders are also carried by dataclass as shown in L23–25 of Listing 1. At least, all decoders always return not only the output texts but also expected scores and their indices in the list of hypotheses. This allows us to analyze where the output text came from when the hypothesis set is constructed from the multiple generation systems. In addition, the decoders can return the reranked N-best lists; thus, it can show the N-best output texts and their expected scores.
3.5 Extensibility
Metrics and decoders are easily customized using the abstract classes. By inheriting these classes, a new class can leverage the predefined common methods and needs only to implement specific methods minimally, e.g., .score() for metrics. If the required methods are not implemented, an exception error will be raised. The necessary methods can be implemented according to the type annotations and docstrings of the parent class.
4 Experiments
Decoding Utility Est. bleu chrf comet bleurt kiwi
MAPϵ 23.8 50.1 75.6 54.9 74.2
[…]
QE kiwi 22.5 51.8 82.1 58.4 83.6
MBR bleu MC 25.6 52.9 75.2 54.8 73.7
MB 24.2 50.5 75.5 54.8 74.1
chrf MC 24.4 54.2 76.3 56.4 75.1
MB 24.4 51.2 76.0 55.7 74.8
comet MC 24.1 52.8 83.9 58.7 79.5
MB 24.1 51.1 81.0 57.1 77.5
bleurt MC OOT OOT OOT OOT OOT
MB OOT OOT OOT OOT OOT
RAMBR bleu MC 25.6 52.9 75.2 54.9 73.8
MB 24.2 50.5 75.5 54.8 74.1
chrf MC 24.3 54.2 76.1 56.3 75.0
MB 24.4 51.1 75.8 55.6 74.7
comet MC 23.9 52.9 83.0 58.1 79.1
MB 24.3 51.0 80.1 56.6 77.0
CBMBR comet MC 23.2 52.4 83.8 58.3 79.2
MB 22.6 51.6 82.2 57.4 77.8
Prune- bleu MC 25.7 52.9 75.0 54.6 73.6
MBR MB 24.3 52.0 75.5 54.9 74.2
chrf MC 24.5 54.3 76.2 56.3 75.1
MB 24.6 52.8 76.1 55.9 75.0
comet MC 24.0 52.8 83.9 58.7 79.5
MB 24.4 52.5 82.1 57.9 78.4
bleurt MC 22.7 51.7 77.6 63.4 75.6
MB 24.1 52.2 77.3 59.4 75.9
PMBR bleu MC 25.0 52.2 74.2 54.4 72.8
MB 24.9 52.0 73.7 54.2 72.3
chrf MC 23.0 52.9 74.3 55.0 73.1
MB 22.9 52.7 74.1 54.8 72.8
comet MC 23.9 52.7 83.8 58.7 79.5
MB 23.6 51.8 82.6 57.7 78.6
bleurt MC 22.7 51.8 77.7 63.3 75.7
MB 23.1 51.5 77.2 60.8 75.8
Oracle bleu 44.4 62.8 76.1 59.5 71.6
chrf 39.5 66.6 77.7 62.0 72.9
comet 30.2 58.0 87.0 64.1 80.4
bleurt 31.0 58.2 79.7 71.0 75.0
Table 2: Experimental results on the WMT’22 En–De general MT task.
4.1 Setup
Using our mbrs, we conducted translation experiments on WMT’22 general translation task (Kocmi et al., 2022) in En–De, and evaluated the translation quality and speed. We compared various MBR decoding methods: vanilla (Eikema and Aziz, 2020, 2022), reference aggregation (DeNero et al., 2009; Vamvas and Sennrich, 2024), centroid-based aggregation (Deguchi et al., 2024), confidence-based pruning (Cheng and Vlachos, 2023), and probabilistic MBR (Trabelsi et al., 2024). In addition to MBR decoding, we compared other decoding methods, MAP decoding (MAP) and N-best reranking using a quality estimation model (QE). We also measured the upper bound of the translation quality (Oracle) by selecting translations that maximize the evaluation metric using the reference translations.
[…]
We evaluated the translation quality on bleu (Papineni et al., 2002), chrf (Popović, 2015), comet (Rei et al., 2022a), bleurt (Sellam et al., 2020), and COMETkiwi (kiwi) (Rei et al., 2022b). We employed kiwi for QE reranking, and other metrics for the utility functions of MBR decoding and the oracle. The detailed setup is described in Appendix D.
4.2 Results
Step MBR PruneMBR RAMBR PMBR
Prune 950.7
Aggregate 494.3
ALS 159.3
Utility 15789.8 12158.0 104.3 2268.6
E2E 16076.0 11206.8 605.6 2704.2
Table 3: The execution time when bleu is employed as the utility function.
Prune- RA- CB-
Step QE MBR MBR MBR MBR PMBR
Rerank 304.0
Encode 316.2 310.4 325.0 324.0 319.2
Prune 5.6
Aggregate 0.0
Clustering 21.2
ALS 5.9
Utility 277.5 45.3 0.3 11.4 21.4
E2E 522.3 618.2 361.6 325.9 363.9 567.3
Table 4: The execution time when comet is employed as the utility function.
Quality of decoded texts
[…]
To summarize, mbrs can compare various metrics and decoding methods systematically.
Execution time
Table 3 and 4 show the execution time in milliseconds (msec) per sentence when we used bleu and comet as the utility functions, respectively. “QE” in Table 4 shows the reference time of QE reranking with kiwi. In the tables, “Step” indicates each step of decoding, and “E2E” reports the end-to-end total time including miscellaneous processes. Both tables demonstrate that alternative algorithms of the vanilla MBR decoding improved the speed. The additional results with other utility functions are shown in Appendix E.
mbrs can measure the time spent for each profiling code block as shown in the tables, it is helpful for identifying bottlenecks and developing new algorithms.
Distribution of the expected utility
Figure 2: The distribution of expected utility scores in the set of hypotheses. The left and right ones show the examples that have the maximum and minimum variance of the expected utility in the test set, respectively.
Figure 3: The empirical cumulative distribution of ranks selected by MBR decoding from descending rankings of output probabilities.
As described in Section 3.4, mbrs can present the additional information of the output texts. We visualized the distribution of expected utility scores in the hypothesis set by using them. Figure 2 compares the distribution of the expected utility scores between the test cases that have the maximum and minimum variance in the test set. We observed that there are cases in which the expected utility varies significantly depending on each hypothesis and cases in which it does not vary much. The latter result suggests that the N-best outputs of MBR decoding are sometimes similar to each other.
mbrs enables us this analysis by returning the metadata of the output texts.
Visualize the decision of MBR decoding
mbrs also returns which indices in the hypotheses are selected by MBR decoding. Figure 3 shows the empirical cumulative distribution of ranks selected by MBR decoding from descending rankings of output probabilities. In the figure, the x-axis indicates the 0-indexed rank of output probabilities in the hypothesis set, i.e., 0 is the highest probability in the hypotheses set, and the y-axis indicates the ratio of the number of selected by MBR decoding when using COMET as the utility function.
From the results, by using mbrs, we could observe that MBR decoding does not always select high-probability texts, i.e., high-quality texts and high-probability texts are different, as mentioned by Freitag et al. (2022).
5 Conclusion
This paper describes mbrs, a library for MBR decoding, which implements various metrics and decoding methods. It is designed to ensure transparency, reproducibility, and extensibility for researchers and developers. We will continue to implement the latest metrics and decoding methods. We hope our mbrs will be used as a tool to further improve the quality of text generation models.
Ethics Statement and Broader Impact
MBR decoding selects output texts from a set of a candidate list generated by text generation models; hence, if the systems generate toxic text, it may be selected depending on the utility functions. In addition, if a utility function is biased and not aligned with human preferences, MBR decoding is more likely to generate biased texts. Since MBR decoding reflects human preferences in the output texts through the utility function, it must be carefully designed to avoid generating harmful texts.
Appendix A Limitations
We measured the execution times only in a single run; thus, the speed may differ when different computer architectures are used.
Currently, our repository mainly implements metrics for translation tasks. Implementing evaluation metrics other than the translation task is a future work.
Appendix B License
mbrs is published under the MIT license. Our implementation calls the COMET and BLEURT models, licensed under the Apache-2.0 license, and the COMETkiwi and XCOMET models, licensed under the CC BY-NC-SA 4.0 license. We used the test sets of WMT’22 general translation tasks published under the following policy: “The data released for the WMT General MT task can be freely used for research purposes”.
Appendix C Links to Our Project
Appendix D Detailed Experimental Setup
In CBMBR, we set the number of centroids to 64. The reduction factor of PMBR was set to 8 and we set $\alpha=0.99$ in PruneMBR.
We measured the speeds on an NVIDIA RTX 6000 Ada GPU with a batch size of 256 and the half-precision computation for comet, kiwi, and bleurt metrics, and on the 32 core CPUs [^10] Intel (R) Xeon (R) Platinum 8160 for bleu and chrf metrics.
Table 5 lists the model names that we used for the evaluation metrics.
Metric Model
comet Unbabel/wmt22-comet-da
kiwi Unbabel/wmt22-cometkiwi-da
bleurt BLEURT-20-D12
Table 5: Model list of evaluation metrics that we used in the experiments.
Appendix E Additional Experiments
E.1 Execution time on the other metrics
Step MBR PruneMBR RAMBR PMBR
Prune 958.4
Aggregate 1090.8
ALS 96.2
Utility 36540.7 10792.9 539.7 4694.3
E2E 36695.3 11751.7 1642.5 5019.6
Table 6: The execution time when chrf is employed as the utility function.
Step MBR PruneMBR PMBR
Prune 322.0
Aggregate
ALS 5.5
Utility OOT 3026.8 38647.1
E2E OOT 3349.0 38797.8
Table 7: The execution time when bleurt is employed as the utility function.
Table 6 and 7 show the execution times of MBR decoding on the WMT’22 En–De translation task when chrf and bleurt are used as the utility functions, respectively.
E.2 Additional translation experiments on other language directions
[…]
De–En En–Zh Zh–En En–Ja Ja–En
Decoding Utility Est. bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi
MAPϵ 26.0 50.8 78.4 62.0 75.4 25.1 25.0 76.4 53.0 73.3 12.1 35.0 67.2 49.5 65.4 15.1 22.6 79.2 46.1 77.4 9.0 29.2 68.0 46.5 68.5
[…]
QE kiwi 23.8 50.7 82.2 63.3 81.9 25.8 25.1 83.0 55.7 82.8 15.4 46.6 76.9 55.0 77.7 16.3 26.1 86.4 50.8 86.7 10.8 37.0 76.3 52.3 80.1
MBR bleu MC 27.2 52.1 78.5 62.0 75.3 29.8 27.8 77.3 53.8 74.2 17.2 46.8 71.9 51.8 69.4 17.9 26.2 81.5 48.2 80.0 11.7 35.5 70.1 47.4 69.9
MB 26.4 51.1 78.5 61.9 75.3 25.6 25.5 76.5 53.2 73.4 12.4 35.4 67.4 49.7 65.7 15.6 23.1 79.9 46.8 78.0 9.4 29.6 68.1 46.3 68.5
chrf MC 27.5 53.4 79.0 62.8 76.0 29.5 28.5 77.6 53.8 74.4 17.3 48.9 72.8 53.1 70.4 16.8 26.7 81.9 49.1 80.9 12.0 38.4 70.9 49.5 71.1
MB 26.8 51.9 78.8 62.5 76.0 25.4 25.8 76.8 53.4 73.7 12.3 35.9 67.9 50.5 66.2 15.6 23.4 80.1 47.3 78.6 9.8 31.0 68.8 47.7 69.5
comet MC 25.7 51.9 83.0 63.7 78.5 28.1 26.5 84.1 55.4 78.7 15.8 46.7 77.3 54.1 73.0 16.6 26.0 87.9 50.5 83.5 10.5 36.0 76.7 50.4 74.0
MB 26.2 51.4 81.6 63.2 77.6 25.5 25.7 81.4 54.5 76.6 12.3 35.6 69.9 50.8 67.4 15.7 23.6 84.0 48.7 80.5 9.6 30.4 72.2 48.4 71.2
bleurt MC OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
MB OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
RAMBR bleu MC 27.3 52.1 78.6 62.1 75.4 29.8 27.7 77.3 53.7 74.2 17.2 46.8 72.1 52.0 69.6 17.7 26.0 81.5 48.0 80.0 11.8 35.6 70.1 47.8 70.0
MB 26.4 51.1 78.5 62.0 75.4 25.4 25.4 76.4 53.2 73.3 12.0 35.3 67.3 49.8 65.6 15.5 23.0 79.7 46.6 77.9 9.4 29.9 68.1 46.7 68.7
chrf MC 27.4 53.4 79.0 62.9 76.0 29.4 28.6 77.5 53.5 74.4 17.1 48.7 72.8 53.1 70.4 16.8 26.8 81.8 49.1 80.8 11.8 38.4 71.0 49.5 71.2
MB 26.9 51.9 78.9 62.6 76.0 25.1 25.7 76.6 53.0 73.6 12.2 35.8 67.8 50.5 66.2 15.3 23.4 80.0 47.3 78.6 9.8 31.1 68.8 47.9 69.5
comet MC 26.0 51.9 82.3 63.5 78.0 28.7 26.9 83.2 55.4 78.0 16.0 46.9 76.6 53.9 72.6 16.9 26.2 87.3 50.5 83.4 10.8 36.2 75.8 50.1 73.4
MB 26.2 51.4 81.0 63.1 77.2 25.6 25.7 80.5 54.3 76.0 12.3 35.7 69.6 50.7 67.2 15.9 23.6 83.3 48.3 80.1 9.4 30.3 71.5 48.1 70.6
CBMBR comet MC 25.1 51.4 83.0 63.3 78.3 27.6 26.0 84.0 55.2 78.5 15.7 46.5 77.2 53.7 72.8 16.5 26.0 87.8 50.1 83.3 10.0 35.4 76.6 50.1 73.5
MB 24.2 50.4 81.9 62.6 77.4 26.9 25.6 82.8 54.4 77.5 14.5 44.5 75.0 52.6 71.1 15.7 24.9 86.4 49.0 82.3 9.3 33.8 74.5 48.8 72.1
Prune- bleu MC 27.1 52.0 78.5 61.9 75.3 29.8 27.8 77.3 53.8 74.2 17.2 46.8 72.0 51.8 69.4 17.9 26.2 81.5 48.2 80.0 11.7 35.4 70.0 47.4 69.8
MBR MB 26.2 51.4 78.6 61.9 75.4 28.5 27.1 77.3 53.7 74.1 15.5 44.1 71.2 51.4 68.9 16.5 25.1 81.3 47.5 79.4 10.1 33.1 69.4 46.9 69.3
chrf MC 27.4 53.5 78.9 62.8 75.9 29.5 28.5 77.6 53.8 74.5 17.2 48.8 72.8 53.2 70.5 16.9 26.8 81.8 49.1 80.9 11.9 38.3 70.8 49.5 71.0
MB 26.8 52.3 78.8 62.4 75.9 28.8 27.4 77.6 53.9 74.4 15.9 45.2 71.7 52.2 69.4 16.6 25.4 81.5 48.0 79.9 10.8 34.9 70.2 48.4 70.3
comet MC 25.8 51.9 83.0 63.7 78.4 28.1 26.5 84.1 55.5 78.6 15.9 46.8 77.3 54.1 72.9 16.5 26.0 87.9 50.4 83.6 10.4 35.9 76.6 50.3 73.9
MB 26.3 51.9 81.6 63.3 77.6 28.9 27.3 82.3 55.1 77.4 15.2 44.3 74.4 52.7 70.9 17.0 25.7 86.1 49.7 82.3 10.0 34.1 73.9 49.2 72.2
bleurt MC 24.9 51.3 79.4 65.1 76.4 27.3 25.9 78.0 58.7 75.0 15.7 46.5 73.6 56.7 71.3 15.8 25.5 82.5 53.9 80.6 10.2 35.2 71.3 53.2 71.4
MB 26.0 51.5 79.1 63.5 76.1 28.3 26.9 77.7 55.7 74.6 15.1 44.2 72.0 53.7 69.7 16.2 25.0 81.9 50.0 80.0 9.9 33.2 70.4 50.0 70.4
PMBR bleu MC 26.4 51.4 78.2 61.8 75.0 29.0 27.1 77.0 53.7 73.9 16.9 46.5 71.8 51.7 69.2 17.4 25.7 81.1 48.1 79.7 11.2 34.9 69.9 47.4 69.8
MB 26.4 51.5 78.0 61.6 74.7 28.9 27.0 76.7 53.2 73.7 16.5 45.8 71.4 51.4 68.8 17.3 25.7 81.2 47.8 79.6 11.1 34.6 69.4 47.1 69.2
chrf MC 26.0 52.2 78.3 62.1 75.1 29.0 27.9 77.3 53.7 74.2 16.0 47.6 72.2 52.5 69.8 16.4 26.1 81.3 48.5 80.2 11.1 37.2 70.4 49.0 70.6
MB 25.7 52.1 78.1 61.9 75.0 28.7 27.8 76.9 53.1 73.8 15.9 47.4 72.0 52.4 69.6 16.3 26.0 81.1 48.2 79.9 11.2 37.2 70.1 48.7 70.2
comet MC 25.9 51.9 82.8 63.6 78.3 28.3 26.6 83.9 55.5 78.5 15.9 46.7 77.1 54.0 72.9 16.7 26.0 87.7 50.3 83.4 10.5 35.8 76.5 50.3 73.8
MB 25.7 51.5 82.1 63.5 77.9 27.0 26.1 82.5 54.9 77.4 14.1 41.5 73.5 52.3 70.0 16.0 24.7 85.9 49.4 82.1 10.0 32.6 74.0 49.1 72.1
bleurt MC 25.0 51.4 79.4 65.1 76.4 27.6 26.2 78.0 58.6 74.9 15.5 46.4 73.5 56.6 71.2 15.8 25.5 82.5 53.7 80.6 10.1 35.2 71.5 53.2 71.6
MB 25.5 51.4 79.2 64.1 76.3 27.7 26.3 77.6 56.9 74.5 14.7 44.1 72.0 54.4 69.6 16.0 24.7 82.0 51.2 79.9 9.6 33.5 70.5 51.1 70.7
Oracle bleu 46.4 63.4 80.9 66.6 74.3 45.5 41.9 80.9 61.4 74.2 31.9 54.9 74.4 56.3 69.0 34.7 39.0 83.4 54.4 78.7 28.0 46.2 73.2 53.1 69.5
chrf 43.5 66.0 81.3 67.4 74.8 45.6 42.8 81.0 61.3 74.1 28.2 58.2 75.5 57.8 70.0 33.2 41.4 84.2 56.3 79.2 24.0 50.4 74.5 55.2 70.8
comet 34.4 58.3 86.4 68.5 78.9 35.4 33.4 87.6 61.8 79.3 20.5 50.8 81.2 58.9 73.9 22.4 31.6 90.6 56.8 83.6 15.9 41.2 81.9 56.3 75.5
bleurt 35.7 59.0 82.5 72.6 76.3 36.3 34.2 81.9 68.6 75.6 21.3 50.9 76.6 63.8 71.6 23.5 33.0 85.3 64.3 80.4 16.3 41.6 75.9 62.2 72.8
[…]